A body of mass 6kg rest on an inclined plane.the normal reaction is R and the limiting frictional force is f.if f is 30N and g is 10ms-².what is the angle of inclination.
1
Expert's answer
2018-11-13T12:57:10-0500

The Newton’s second law gives

R+f+mg=0

In projection on the x-axis

f-mg sin⁡θ=0

So

sin⁡θ=f/mg=30/(6×10)=1/2

θ=30°